Molly and Mike are mowing the lawn. When Mike mows the lawn by himself it takes him 3 hours, but with Molly helping it takes 2 hours. How long does it take Molly to mow the law by herself?
Answer by richwmiller(17219)

2/m+2/3=1
2/m=1/3
m=6